# Enhance Mobile Agents Thinking Process Via Iterative Preference Learning

### v2

@@ 1 Introduction
VLM-based mobile agents (wang2023enabling, ; ding2024mobileagent, ) have attracted considerable attention due to their ability to seamlessly interact with mobile graphical user interfaces (GUIs) and their potential to autonomously perform daily tasks. Since actions are not directly specified in user instructions, mobile agents benefit from generating intermediate thoughts aligned with the current GUI context. Recent work such as AITZ(zhang2024android, ) has demonstrated that the Chain of Action-Planning Thoughts (CoaT) pattern—resembling the slow-thinking “System 2” process—is particularly effective in GUI domains.
@@
However, directly applying supervised fine-tuning (SFT) on CoaT trajectories may cause overfitting, leading the model to be trapped in fixed reasoning patterns. To address this limitation, recent studies in the general domain have explored self-training strategies. These approaches typically utilize the correctness of the final answer in output as a reward signal to train the model(luong2024reft, ). While effective in some contexts, relying solely on final answers overlooks the quality of intermediate reasoning steps, which can result in reward hacking and suboptimal reasoning processes. Some search-based approaches, such as ReST-MCTS (xie2024monte, ), tackle this problem by learning a process reward model (PRM) to evaluate individual reasoning steps. However, these approaches often require large-scale manual annotation of intermediate steps (guo2025deepseek-r1, ). This challenge is especially severe in the Mobile GUI Agent domain. Unlike text-based tasks in coding or math, GUI environments rely on real devices or simulators, making step-level reward annotation significantly more costly and labor-intensive.
@@
To address these limitations, we propose an interactive sampling framework that constructs a CoaT-tree based on Monte Carlo Tree Search (MCTS). Rather than relying solely on final answers, our method assigns scores to individual reasoning steps, enabling the construction of thinking-level Direct Preference Optimization (DPO) pairs. Specifically, we perform multi-turn dialogue with a vision-language model (VLM) to incrementally build a CoaT-tree, where each node corresponds to a sampled response at a given reasoning step, conditioned on the dialogue history. This hierarchical structure captures diverse reasoning paths and facilitates fine-grained assessment of intermediate thoughts. We first assign rewards to the leaf node, and then propagate these signals backward through the CoaT-tree to earlier reasoning steps. Based on the resulting values, we construct thinking-level DPO pairs to help agents optimize both final actions and the overall quality of their reasoning.
@@
To mitigate the lack of diversity after warm-up SFT, we adopt an instruction evolution strategy. Specifically, we generate diverse Q&A pairs grounded in real mobile UI screenshots from downstream datasets. These Q&A pairs serve two purposes: (1) prevent agents from overfitting to static downstream instructions by introducing varied reasoning contexts, and (2) improve agents’ understanding of UI layouts through visually grounded question-answering. We evaluate our approach on the CoaT dataset AITZ and long-horizon dataset AMEX, where it outperforms state-of-the-art GUI-agent continual pretraining agents such as OS-ATLAS (wu2024atlas, ) (+4.04%) and UI-TARS (qin2025ui, ) (+3.54%). Furthermore, experiments on the AndroidControl dataset demonstrate the strong generalization capability of our method to unseen apps and instructions (tasks). Under limited training resources, IPL consistently outperforms naive-DPO using only half of the data for one iterative training round (+4.5%), or one-fifth of the data for two iterative training rounds (+0.3%). Analytical experiments show instruction evolution simultaneously improves both the diversity and quality of reasoning.
@@ 2.1 Mobile GUI Agent
LLMs (achiam2023gpt, ) are increasingly used as autonomous agents for mobile interaction (lieffects, ; wen2023empowering, ). With the rapid development of vision-language models (VLMs), researchers build mobile GUI agents (yang2023appagent, ; zheng2024gpt, ; qin2025ui, ; teamqwen2, ) and multi-agent frameworks (ding2024mobileagent, ; li2024appagent, ; wang2024mobile, ) based on closed-source VLMs. Meanwhile, some researchers focus on training agents with stronger element grounding (cheng2024seeclick, ; wu2024atlas, ), page navigation (niu2024screenagent, ; lu2024gui, ; gou2024navigating, ; wang2025fedmobileagent, ), GUI understanding (you2024ferret, ; baechler2024screenai, ) and task planning capabilities (zhang2024llamatouch, ; nong2024mobileflow, ; xu2024aguvis, ; qinghong2024showui, ; dorka2024training, ) based on open-source VLMs. Our method organizes trajectory data into multi-rounds of dialogues based on the CoaT thinking pattern, preventing the agent becomes an action model with limited capabilities.
@@ 2.2 Reinforcement Learning
The algorithms applied in natural language processing to align with human preferences include Direct Preference Optimization (DPO) (rafailov2024direct, ), Identity Preference Optimization (IPO) (azar2024general, ), Kahneman-Tversky Optimization (KTO) (ethayarajh2023human, ), and Proximal Policy Optimization (PPO) (schulman2017proximal, ). Specifically, ReFT (luong2024reft, ) adopts reinforcement learning as a fine-tuning paradigm to improve performance on math problems. ReST-MCTS* (zhang2024rest, ) focuses on the higher-quality step reward, where the process reward model is important. Xie, et al. (xie2024monte, ) label the preference via MCTS based on feedback from self-evaluation. For mobile GUI agents, Digirl (bai2024digirl, ) and Distrl (wang2024distrl, ) use online trajectory collection to improve the generalization of agents whose process is very slow. Reachagent (wu2025reachagent, ) uses DPO training to compare the quality of multiple actions. Our method uses T-DPO to optimize the thinking process, while the step value is calculated directly by rules without unstable PRMs.
Figure 1: Overview of iterative preference learning framework. The left part presents the process of warm-up fine-tuning a general VLM to a mobile GUI domain agent with basic capabilities. The mid and right parts represent the iterative CoaT thinking-level sampling and DPO self-training process.
@@ 3 Methodology
In this section, we first introduce the multi-round thinking process formulation  (§ 3.1) and explain our method. As shown in Figure 1, our method starts with instruction evolution strategy (§ 3.2) to enhance output diversity after warm-up SFT. Then, a CoaT-tree through iterative sampling (§ 3.3) is employed for each action. Every leaf node represents a complete action and is scored using a rule-based reward function. We then backpropagate the rewards along the tree to assign credit to intermediate reasoning steps. This process yields thinking-level contrastive pairs for DPO, which further improves the model’s reasoning ability. The detailed process is presented in Algorithm 1.
3.1 Multi-round Thinking Process Formulation
@@
We formulate action $\hat{a}_{i}$ in the CoaT reasoning process as a multi-round dialogue $\hat{a}_{i}=[s_{1},s_{2},s_{3},s_{4}]$, where $s_{i}$ represents description, action-thought, action-decision, and grounding, respectively. This thinking paradigm based on the thinking–decision–grounding triplet, has been widely validated as effective in previous GUI works (shen2024falcon, ; zhang2024android, ; qin2025ui, ; cheng2024seeclick, ). So the reasoning process can be formulated as:
@@
$P$ represents each round of dialogue input prompt, $I$ is the task instruction, $u$ is the current GUI, and $\hat{a}_{i}$ is the step $i$ history action. Agents perform poorly when decoding the entire reasoning process in a single step, which is because image modal $u$ dominates the input tokens, surpassing textual instructions $I$ and action history $\hat{a}_{i}$, and diverting their attention away from the textual details. During autoregressive training, the agent is unaware that producing a final answer conforming to the required format is indispensable throughout the reasoning process. Multi-round thinking process effectively mitigates this problem, because additional dialogue steps guarantee a final answer is generated:
@@ 3.2 Instruction Evolution
As introduced in the previous section, the CoaT patterns in the mobile agent domain are often fixed. Agents tend to overfit on these fixed paradigms and struggle to generate diverse reasoning (in Sec 4.4). To address this challenge, we augment the original training trajectories $\mathcal{T}$ by attaching additional Q&A annotations to UI screenshots through an instruction evolution process, resulting in a new dataset $\mathcal{Q}$ with more diverse instruction forms. Specifically, as shown in Figure 2, the evolution process consists of three levels:
Level I: General GUI Q&A tasks. Grounding, Reference (Ref), and Page Descriptions are aimed at enhancing the agent’s foundational capabilities. These tasks (liu2024autoglm, ; yang2024aria, ) are proven to be the core capabilities of GUI agents during the pre-training.
Level II: Widget caption and relationships. Descriptions of widget functions and the nested partition relationships between widgets. These tasks help agents understand the relationships between widgets, as previous work (deng2024mobile, ) has found that agents tend to click on the textview, even in scenarios where the textview and the button are separate.
Level III: GUI advanced FAQ. The advanced FAQ includes more complex Q&A, such as descriptions of part of the structural framework of the page, expectations and predictions about page navigation triggered by control interactions.
Warm-up Supervised Fine-tuning: To develop agents with standard thinking format and expand the reasoning space, we mix $\mathcal{T}$ and the instruction evolution data $\mathcal{Q}$, then perform warm-up SFT on $\mathcal{D}=\Big{\{}\mathcal{T},\mathcal{Q}\Big{\}}=\Big{\{}(u,e)^{(i)}\Big{\}}_{i=1}^{|\mathcal{D}|}$, where $u$ represents the prior knowledge (instructions, screenshot and action history) from $\mathcal{T}$ or the questions from $\mathcal{Q}$. $e$ is the reasoning process from $\mathcal{T}$ or the answer from $\mathcal{Q}$ which is organized into multi-round dialogues. To ensure output diversity, we select an earlier checkpoint with better potential correct space and diverse output to serve as the seed policy model. More details can be seen in Appendix B.
@@ 3.3 Iterative Preference Learning
After warm-up SFT, the agent acquires basic GUI capabilities. To further enrich the original thinking process, we construct a CoaT-tree through iterative sampling for each reasoning step. Based on rule-based rewards, thinking-level DPO pairs are filtered to optimize agents’ reasoning process.
Iterative Sampling & Rule-based Reward. We iteratively sample each reasoning step along the CoaT paradigm (zhang2024android, ). The $\mathcal{K}$ sampling results $(\hat{s}_{t}|\hat{s}_{1:t-1})^{\mathcal{K}}$ at step $t$ can be expressed as:
@@
$\small v(s_{t})=\begin{cases}1,&d(s_{t},a^{*})_{[x,y]}<=d_{min}\vee dir(s_{t}\sim a^{*})_{scroll}\\ 1-(1-v_{format})\cdot\frac{d(s_{t},a^{*})_{[x,y]}}{d_{max}-d_{min}},&\\ &d(s_{t},a^{*})_{[x,y]}>d_{min}\\ v_{type}+v_{format}+(1-v_{type}-v_{format})\cdot F_{1},&F_{1}(s_{t},a^{*})_{text}\\ v_{type},&type(s_{t}\sim a^{*})\\ 0,&else\end{cases}$ (7)
The reward score $v(s_{t})$ ranges from 0 to 1, with a fully correct prediction receiving a score of 1. We use $v_{\text{type}}$ to indicate whether the predicted action type matches the ground truth.
For click and Input actions, we further evaluate their internal structure. Specifically, we introduce $v_{\text{format}}$ to indicate whether essential arguments (e.g., coordinates or text) are correctly extracted. The final reward is computed based on the similarity between the prediction and the ground truth:
  • Click: A distance-based score between the predicted and ground-truth coordinates, scaled between $d_{\min}$ and $d_{\max}$.
  • Input: The $F_{1}$ score between the predicted and ground-truth text.
The full reward function is defined in Equation 7 and further discussed in Section C. Based on the structure of the CoaT-tree, we recursively compute the value of each intermediate reasoning step. Specifically, the value of $s_{t-1}$ is computed as the average value of its $\mathcal{K}$ sampled continuations at $s_{t}$:
@@
Contrastive Data Filter. After obtaining the sampling tree and node values, we evaluate the quality of the trees and extract contrastive data. The quality of the sampling trees are =$\mathcal{R}=\{\alpha,\beta,\gamma\}$, and the number of $\alpha,\beta,\gamma$ are calculated as follows:
$\small{\alpha=\frac{\Big{|}\big{\{}\mathcal{S}^{(i)}\mid\forall v_{k}\in\mathcal{S}^{(i)},v_{k}=1\big{\}}\Big{|}}{\sum_{i=1}^{|\mathcal{T}|}|(u,e)^{(i)}|}}$ (9)
$\small{\beta=\frac{\Big{|}\big{\{}\mathcal{S}^{(i)}\mid\exists v_{k},v_{k^{\prime}}\in\mathcal{S}^{(i)},v_{k}=1,v_{k^{\prime}}\neq 1\big{\}}\Big{|}}{\sum_{i=1}^{|\mathcal{T}|}|(u,e)^{(i)}|}}$ (10)
$\small{\gamma=\frac{\Big{|}\big{\{}\mathcal{S}^{(i)}\mid\forall v_{k}\in\mathcal{S}^{(i)},v_{k}\neq 1\big{\}}\Big{|}}{\sum_{i=1}^{|\mathcal{T}|}|(u,e)^{(i)}|}}$ (11)
@@
0:  base VLM $\pi$, advanced annotated model $R_{SoTA}$, step-level trajectory data $\mathcal{T}$, instruction evolution Q&A set $\mathcal{Q}$, number of sampling $\mathcal{K}$, golden action $a^{*}$, value function $v$, the sampled CoaT data $D$, number of iterations $\mathcal{N}$. 1:  for $i=1$ to $N_{0}$ do 2:     $\mathcal{Q}^{*}\leftarrow$ instruction_evolution($R_{\text{SoTA}},\mathcal{T}$) // instruction evolution by GPT-4o 3:     $\mathcal{Q}\leftarrow$ human_evaluation($h,\mathcal{Q}^{*}$) // human filter 4:  end for 5:  $\pi_{S_{0}}\leftarrow$ Warm-up__\__SFT($\pi,\mathcal{T},\mathcal{Q}$) // fine-tune seed model 6:  for $n=1$ to $\mathcal{N}$ do 7:     for $i=1$ to $|\mathcal{T}|$ do 8:        $D_{i}\leftarrow$ generate__\__sampling__\__thought($\pi_{S_{n-1}}$, $\mathcal{T}_{i}$, $\mathcal{K}$) // CoaT Sampling 9:        $V^{leaf}_{i}\leftarrow v(D_{i},a^{*}_{i})$ // match and calculate leaf values using Eq(7) 10:        $V^{intra}_{i}\leftarrow$ recursive__\__calculate($D_{i},V^{leaf}_{i}$) // recursive intra node values using Eq(8) 11:        $D^{+}_{i},D^{-}_{i}\leftarrow$ contrastive_data_filter($D_{i},V_{i}$) // filter positive and negative data using Eq(12, 13) 12:     end for 13:     $\pi_{S_{n}}\leftarrow$ DPO($\pi_{S_{k-1}},D^{+},D^{-}$) // DPO self-training reference model 14:  end for 14:  $\pi_{S},D_{G},\mathcal{Q}$
@@
Direct Preference Optimizing. After CoaT thinking-level Iterative Sampling, several positive and negative example pairs are collected. During this stage, the agent policy undergoes updates through the DPO loss of data pairs mentioned above (rafailov2024direct, ). To further refine the agent’s performance post-optimization, we employ the updated agent as the new base agent to continue collecting contrastive CoaT-action level pairs for additional DPO training. This iterative process is maintained until the agent reaches the performance bottleneck on the test set.

### v3

# MobileIPL: Enhancing Mobile Agents Thinking Process via Iterative Preference Learning

@@ 1 Introduction
VLM-based mobile agents (Wang et al., 2023; Ding, 2024) have attracted considerable attention due to their ability to seamlessly interact with mobile graphical user interfaces (GUIs) and their potential to autonomously perform daily tasks. Since actions are not directly specified in user instructions, mobile agents benefit from generating intermediate thoughts aligned with the current GUI context. Recent work such as AITZ(Zhang et al., 2024b) has demonstrated that the Chain of Action-Planning Thoughts (CoaT) pattern—resembling the slow-thinking “System 2” process—is particularly effective in GUI domains.
@@
However, directly applying supervised fine-tuning (SFT) on CoaT trajectories may cause overfitting, leading the model to be trapped in fixed reasoning patterns. To address this limitation, recent studies in the general domain have explored self-training strategies. These approaches typically utilize the correctness of the final answer in output as a reward signal to train the model(Luong et al., 2024). While effective in some contexts, relying solely on final answers overlooks the quality of intermediate reasoning steps, which can result in reward hacking and suboptimal reasoning processes. Some search-based approaches, such as ReST-MCTS (Xie et al., 2024), tackle this problem by learning a process reward model (PRM) to evaluate individual reasoning steps. However, these approaches often require large-scale manual annotation of intermediate steps (Guo et al., 2025). This challenge is especially severe in the Mobile GUI Agent domain. Unlike text-based tasks in coding or math, GUI environments rely on real devices or simulators, making step-level reward annotation significantly more costly and labor-intensive.
@@
To address these limitations, we propose an iterative sampling framework that constructs a CoaT-tree based on Monte Carlo Tree Search (MCTS). Instead of relying on a PRM, we score each reasoning step and construct thinking-level DPO (T-DPO) pairs without manual step annotation. Specifically, we perform multi-turn dialogue with a vision-language model (VLM) to incrementally build a CoaT-tree, where each node corresponds to a sampled response at a given reasoning step, conditioned on the dialogue history. This hierarchical structure captures diverse reasoning paths and facilitates fine-grained assessment of intermediate thoughts. We first assign rewards to the leaf node, and then propagate these signals backward through the CoaT-tree to earlier reasoning steps. Based on the resulting values, we construct thinking-level DPO pairs to help agents optimize both final actions and the overall quality of their reasoning.
@@
To mitigate the lack of diversity after warm-up SFT, we adopt an instruction evolution strategy. Specifically, we generate diverse Q&A pairs grounded in real mobile UI screenshots from downstream training datasets. These Q&A pairs serve two purposes: (1) prevent agents from overfitting to static downstream instructions by introducing varied reasoning contexts, and (2) improve agents’ understanding of UI layouts through visually grounded question-answering. We evaluate our approach on the CoaT dataset AITZ and long-horizon dataset AMEX, where it outperforms state-of-the-art GUI-agent continual pretraining agents such as OS-ATLAS (Wu et al., 2024) (+4.04%) and UI-TARS (Qin et al., 2025) (+3.54%). Furthermore, experiments on the AndroidControl dataset demonstrate the strong generalization capability of our method to unseen apps and instructions (tasks). Under limited training resources, IPL consistently outperforms naive DPO using only half of the data for one iterative training round (+4.5%), or one-fifth of the data for two iterative training rounds (+0.3%). Analytical experiments show instruction evolution simultaneously improves both the diversity and quality of reasoning.
@@ 2.1 Mobile GUI Agent
LLMs (Achiam et al., 2023) are increasingly used as autonomous agents for mobile interaction (Li et al., 2024b; Wen et al., 2023). With the rapid development of vision-language models (VLMs), researchers build mobile GUI agents (Yang et al., 2023; Zheng et al., 2024; Qin et al., 2025; Team, ) and multi-agent frameworks (Ding, 2024; Li et al., 2024c; Wang et al., 2024a) based on closed-source VLMs. Meanwhile, some researchers focus on training agents with stronger element grounding (Cheng et al., 2024; Wu et al., 2024), page navigation (Niu et al., 2024; Lu et al., 2024; Gou et al., 2024; Wang et al., 2025), GUI understanding (You et al., 2024; Baechler et al., 2024) and task planning capabilities (Zhang et al., 2024c; Nong et al., 2024; Xu et al., 2024; Qinghong Lin et al., 2024; Dorka et al., 2024) based on open-source VLMs. Our method organizes trajectory data into multi-turns of dialogues based on the CoaT thinking pattern, preventing the agent becomes an action model with limited capabilities.
@@ 2.2 Reinforcement Learning
The algorithms applied in natural language processing to align with human preferences include Direct Preference Optimization (DPO) (Rafailov et al., 2023), Identity Preference Optimization (IPO) (Azar et al., 2024), Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2023), and Proximal Policy Optimization (PPO) (Schulman et al., 2017). Specifically, ReFT (Luong et al., 2024) adopts reinforcement learning as a fine-tuning paradigm to improve performance on math problems. ReST-MCTS* (Zhang et al., 2024a) focuses on the higher-quality step reward, where the process reward model is important. Xie, et al. (Xie et al., 2024) label the preference via MCTS based on feedback from self-evaluation. For mobile GUI agents, Digirl (Bai et al., 2024) and Distrl (Wang et al., 2024c) use online trajectory collection to improve the generalization of agents whose process is very slow. Reachagent (Wu et al., 2025) uses DPO training to compare the quality of multiple actions. Our method uses T-DPO to optimize the thinking process, while the step value is calculated directly by rules without unstable PRMs.
Figure 1: Overview of iterative preference learning framework. The left part presents the process of warm-up fine-tuning a general VLM to a mobile GUI domain agent with basic capabilities. The mid and right parts represent the iterative CoaT thinking-level sampling and T-DPO training process.
@@ 3 Methodology
In this section, we first introduce the multi-turn thinking process formulation  (§ 3.1) and explain our method. As shown in Figure 1, our method starts with instruction evolution strategy (§ 3.2) to enhance output diversity in warm-up SFT stage. Then, a CoaT-tree through iterative sampling (§ 3.3) is employed for each action. Every leaf node represents a complete action and is scored using a rule-based reward function. We then backpropagate the rewards along the tree to assign credit to intermediate reasoning steps. This process yields thinking-level contrastive pairs for DPO, which further improves the model’s reasoning ability. The detailed process is presented in Algorithm 1.
3.1 Multi-turn Thinking Process Formulation
@@
We formulate action $\hat{a}_{i}$ in the CoaT reasoning process as a multi-turn dialogue $\hat{a}_{i}=[s_{1},s_{2},s_{3},s_{4}]$, where $s_{i}$ represents description, action-thought, action-decision, and grounding, respectively. This thinking paradigm based on the thinking–decision–grounding triplet, has been widely validated as effective in previous GUI works (Shen et al., 2024; Zhang et al., 2024b; Qin et al., 2025; Cheng et al., 2024). So the reasoning process can be formulated as:
@@
$P$ represents each round of dialogue input prompt, $I$ is the task instruction, $u$ is the current GUI, and $\hat{a}_{i}$ is the step $i$ history action. Agents perform poorly when decoding the entire reasoning process in a single step, which is because image modal $u$ dominates the input tokens, surpassing textual instructions $I$ and action history $\hat{a}_{i}$, and diverting their attention away from the textual details. During autoregressive training, the agent is unaware that producing a final answer conforming to the required format is indispensable throughout the reasoning process. Multi-turn thinking process effectively mitigates this problem, because additional dialogue steps guarantee a final answer is generated:
@@ 3.2 Instruction Evolution
As discussed in the previous section, the CoaT patterns in the mobile agent domain are typically fixed. As a result, agents tend to overfit these static paradigms and struggle to generate diverse reasoning paths after the warm-up SFT training (as detailed in Sec 4.4). To address this issue, we enhance the original training trajectories, denoted as $\mathcal{T}$, by appending additional Q&A annotations to UI screenshots through an instruction evolution process, thereby creating a new dataset $\mathcal{Q}$ with a broader range of instruction formats. Specifically, as shown in Figure 2, the evolution process consists of three levels:
Level I: General GUI Q&A tasks. Grounding, Reference (Ref), and Page Descriptions are aimed at enhancing the agent’s foundational capabilities. These tasks (Liu et al., 2024; Yang et al., 2024) are proven to be the core capabilities of GUI agents during the pre-training.
Level II: Widget caption and relationships. Descriptions of widget functions and the nested partition relationships between widgets. These tasks help agents understand the relationships between widgets, as previous work (Deng et al., 2024) has found that agents tend to click on the textview, even in scenarios where the textview and the button are separate.
Level III: GUI advanced FAQ. Inspired by Shen et al. (2024), we design an advanced FAQ that features more complex Q&A, including descriptions of the page’s structural framework as well as expectations and predictions about navigation outcomes triggered by control interactions.
Warm-up Supervised Fine-tuning: To develop agents with standard thinking format and expand the reasoning space, we mix $\mathcal{T}$ and the instruction evolution data $\mathcal{Q}$, then perform warm-up SFT on $\mathcal{D}=\Big\{\mathcal{T},\mathcal{Q}\Big\}=\Big\{(u,e)^{(i)}\Big\}_{i=1}^{|\mathcal{D}|}$, where $u$ represents the prior knowledge (instructions, screenshot and action history) from $\mathcal{T}$ or the questions from $\mathcal{Q}$, and $e$ is the reasoning process from $\mathcal{T}$ or the answer from $\mathcal{Q}$ which is organized into multi-turn dialogues. To ensure output diversity, we select an earlier checkpoint with better potential correct space and diverse output to serve as the seed policy model. More details can be seen in Appendix B.
@@ 3.3 Iterative Preference Learning
After the warm-up SFT, the agent acquires basic GUI capabilities. We construct a CoaT-tree by iteratively sampling each reasoning step and then assign a score to the leaf nodes based on a rule-based reward function. Using these scores, we generate thinking-level DPO pairs to optimize the agent’s reasoning process.
Iterative Sampling & Rule-based Reward. We iteratively sample each reasoning step along the CoaT paradigm (Zhang et al., 2024b). The $\mathcal{K}$ sampling results $(\hat{s}_{t}|\hat{s}_{1:t-1})^{\mathcal{K}}$ at step $t$ can be expressed as:
@@
$v(s_{t})=\begin{cases}1,&s_{t}=a*\\ v_{type}+score_{match},&type(s_{t}\sim a^{*})\\ 0,&others\end{cases}$ (7)
[…]
The reward score $v(s_{t})$ ranges from 0 to 1, with a fully correct prediction receiving a score of 1. We use $v_{\text{type}}$ and $v_{\text{format}}$ to indicate whether the predicted action type and output format match the ground truth. For click and input actions, we further evaluate their internal structure using smooth rewards based on spatial distance and text match. The final reward is computed from the similarity between the prediction and the ground truth:
  • Click: A distance-based score between the predicted and ground-truth coordinates, normalized to $[0,1]$; smaller distances yield higher scores.
  • Input: The $F_{1}$ score between the predicted and ground-truth strings; greater textual overlap yields higher scores.
The full reward is defined in Equation 7 and discussed further in Section C.
Based on the structure of the CoaT-tree, we recursively compute the value of each intermediate reasoning step. Specifically, the value of $s_{t-1}$ is computed as the average value of its $\mathcal{K}$ sampled continuations at $s_{t}$:
@@
Contrastive Data Filter. After obtaining the sampling tree and node values, we evaluate the quality of the trees and extract contrastive data. The quality of the sampling trees are $\mathcal{R}=\{\alpha,\beta,\gamma\}$, and the number of $\alpha,\beta,\gamma$ are calculated as follows:
$\alpha=\frac{\Big|\big\{\mathcal{S}^{(i)}\mid\forall v_{k}\in\mathcal{S}^{(i)},v_{k}=1\big\}\Big|}{\sum_{i=1}^{|\mathcal{T}|}|(u,e)^{(i)}|}$} (10)
$\beta=\frac{\Big|\big\{\mathcal{S}^{(i)}\mid\exists v_{k},v_{k^{\prime}}\in\mathcal{S}^{(i)},v_{k}=1,v_{k^{\prime}}\neq 1\big\}\Big|}{\sum_{i=1}^{|\mathcal{T}|}|(u,e)^{(i)}|}$} (11)
$\gamma=\frac{\Big|\big\{\mathcal{S}^{(i)}\mid\forall v_{k}\in\mathcal{S}^{(i)},v_{k}\neq 1\big\}\Big|}{\sum_{i=1}^{|\mathcal{T}|}|(u,e)^{(i)}|}$} (12)
@@
0: base VLM $\pi$, advanced annotated model $R_{SoTA}$, step-level trajectory data $\mathcal{T}$, instruction evolution Q&A set $\mathcal{Q}$, number of sampling $\mathcal{K}$, golden action $a^{*}$, value function $v$, the sampled CoaT data $D$, number of iterations $\mathcal{N}$.
1:for $i=1$ to $N_{0}$ do
2:  $\mathcal{Q}^{*}\leftarrow$ instruction_evolution($R_{\text{SoTA}},\mathcal{T}$) // instruction evolution by GPT-4o
3:  $\mathcal{Q}\leftarrow$ human_evaluation($h,\mathcal{Q}^{*}$) // human filter
4:end for
[…]
6:for $n=1$ to $\mathcal{N}$ do
7:  for $i=1$ to $|\mathcal{T}|$ do
8:   $D_{i}\leftarrow$ generate_\_sampling_\_thought($\pi_{S_{n-1}}$, $\mathcal{T}_{i}$, $\mathcal{K}$) // CoaT Sampling
9:   $V^{leaf}_{i}\leftarrow v(D_{i},a^{*}_{i})$ // match and calculate leaf values using Eq(7)
10:   $V^{intra}_{i}\leftarrow$ recursive_\_calculate($D_{i},V^{leaf}_{i}$) // recursive intra node values using Eq(9)
11:   $D^{+}_{i},D^{-}_{i}\leftarrow$ contrastive_data_filter($D_{i},V_{i}$) // filter positive and negative data using Eq(13, 14)
12:  end for
13:  $\pi_{S_{n}}\leftarrow$ DPO($\pi_{S_{k-1}},D^{+},D^{-}$) // DPO self-training reference model
14:end for
[…]